# Traveling supersolid stripe patterns in spin-orbit-coupled Bose-Einstein condensates

### v1

@@ V.2 Bogoliubov spectrum and sound velocities
The occurrence of an energetic instability in the spectrum of moving stripes is consistent with the behavior observed on the plane-wave side of the phase diagram. In this regime, the roton gap progressively decreases as the Raman coupling $\Omega_{R}$ is lowered towards the transition to the stripe phase [67, 89, 90]. Eventually, the gap closes at a spinodal point slightly below the critical value $\Omega_{\mathrm{cr1}}$. Upon further reduction of $\Omega_{R}$, the Bogoliubov modes near the roton minimum acquire negative frequency. This effect can be seen in the dashed lines of Fig. 4(d1)-(d2), although it appears less pronounced in the comoving frame due to the Doppler shift term $-\bm{v}\cdot\bm{k}$. In contrast, the plane-wave spectrum remains dynamically stable for all values of $\Omega_{R}$. This implies that the dynamical instability of moving stripe patterns must vanish once the limit $s_{p}\to\pm 1$ is reached. Our numerical analysis confirms this expectation: beyond a certain threshold, the dynamically unstable region begins to shrink as $|s_{p}|$ increases, and it disappears entirely when $|s_{p}|=1$. This behavior reflects the trend observed in the sound velocities, as discussed below.
@@
Figure 5: Velocities of spin ($b=1$, red lines) and density ($b=2$, blue lines) sound waves as functions of $|s_{p}|$ for (a) $\hbar\Omega_{R}/E_{R}=1.0$ and (b) $\hbar\Omega_{R}/E_{R}=2.6$. Results for both forward ($c_{b,+}$) and backward ($c_{b,-}$) propagation relative to the momentum direction are shown. For the spin sound velocities, we plot both the real parts (solid lines) and the imaginary parts (dashed lines), with the latter being identical for $c_{1,+}$ and $c_{1,-}$. The interaction parameters are the same as in previous figures: $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$.
@@
We now analyze in more detail the long wavelength behavior of the two gapless excitation bands ($b=1,2$), focusing in particular on the velocities of sound waves propagating parallel ($c_{b,+}$) and antiparallel ($c_{b,-}$) to the momentum direction. For $s_{p}>0$, these velocities characterize the linear dispersions of the gapless bands near the edges of the first Brillouin zone, namely, $\omega_{b,k_{x}}\simeq c_{b,+}k_{x}$ near $k_{x}=0$ and $\omega_{b,k_{x}}\simeq c_{b,-}(2k_{1}-k_{x})$ near $k_{x}=2k_{1}$. The same relations hold for $s_{p}<0$, with the roles of $c_{b,+}$ and $c_{b,-}$ interchanged. As illustrated in Fig. 5, for both the spin ($b=1$) and density ($b=2$) sound modes, the velocities in the direction of the momentum differ from those in the opposite direction, except at $s_{p}=0$, where the two become equal. This symmetry at $s_{p}=0$ reflects the underlying symmetry of the excitation spectrum with respect to the center of the Brillouin zone, as discussed earlier.
@@
The density sound velocities $c_{2,\pm}$ are always real and positive. In the limit $|s_{p}|\to 1$, they approach the expressions $c_{2,\pm}\to c_{\pm x}\mp v_{x}$, where $c_{+x}$ ($c_{-x}$) denotes the speed of density sound waves propagating along (against) the $x$ direction in the plane-wave phase. Note that $c_{+x}\neq c_{-x}$ in the presence of nonzero spin-dependent interactions ($g_{ss}\neq 0$[67].
@@
In contrast, the behavior of the spin sound velocities reveals the onset of instability. As $|s_{p}|$ increases, one of the two velocities $c_{1,\pm}$ becomes negative at a certain point, indicating the emergence of an energetic instability, as previously discussed. For high Raman coupling, this instability is soon followed by the appearance of complex sound velocities: both $c_{1,+}$ and $c_{1,-}$ acquire imaginary parts of equal magnitude. These imaginary parts first grow with $|s_{p}|$, then decrease, and vanish as $|s_{p}|\to 1$. In this limit, the real parts tend towards opposite values, satisfying $c_{1,+}=-c_{1,-}$. This behavior confirms that the emergence of energetic instability is a necessary precursor to ensuring continuity of the excitation spectrum slope in the comoving frame at the points $k_{x}=\pm 2k_{1}^{\mathrm{PW}}$ in the plane-wave phase.
VI Conclusions
@@
In this work, we have investigated the properties of traveling supersolid patterns in a spin-orbit-coupled Bose-Einstein condensate. The motion of the density fringes stems from a population imbalance between the two minima of the single-particle dispersion. Antiferromagnetic spin-dependent interactions create a chemical potential difference which sets the stripe velocity. These moving stripes exhibit asymmetric densities of the two spin components, notable spin polarization, reduced contrast, and higher energy than that of stationary stripes. As momentum increases, these effects intensify, eventually leading to the disappearance of the fringes and the emergence of a uniform plane-wave condensate. This indicates that traveling stripes interpolate between the stationary stripe and plane-wave phases. By analyzing the excitation spectrum, we uncover energetic and dynamical instabilities at high momentum, mainly affecting the spin phonon branch that becomes the mode with the roton minimum in the plane-wave phase. Our findings extend prior work [65, 66] on the slow motion of stripes induced by the release of a weak spin perturbation, where the propagation of fringes manifests the crystal Goldstone mode of the stripe phase.
The results presented in this work may be relevant for ongoing experimental investigations of spin-orbit-coupled Bose gases. Notably, moving stripe patterns have already been observed by Ketterle’s group at MIT [20]. In their experiment, a Raman detuning was employed to generate a chemical potential difference between the two minima of the single-particle spectrum, thereby inducing stripe motion without altering the atomic population balance. In contrast, the configuration explored in this work involves population imbalance between the two momentum components, with the resulting stripe motion driven by spin-dependent interactions. This scenario can be realized experimentally by first preparing the Bose-Einstein condensate in a superposition of two atomic hyperfine states with different weights. Raman coupling can then be adiabatically introduced and ramped up to the desired strength. As in Ref. [20], the resulting moving density modulations could be detected using Bragg scattering techniques.
@@ Appendix A Perturbative method
This appendix presents a detailed methodology employed to derive the perturbative formulas discussed in the main text. The general procedure described in Sec. A.1 closely follows the approach of Ref. [58], with appropriate generalizations to incorporate the additional momentum constraint. The first and second order corrections to the order parameter are subsequently derived in Sec. A.2 and Sec. A.3, respectively.
@@ A.1 General outline of the perturbative approach
To perturbatively solve the time-independent Gross-Pitaevskii equation with a momentum constraint, Eq. (8), we first observe that the stripe order parameter (14) depends solely on the combination $k_{1}x^{\prime}$. This allows us to neglect any derivatives with respect to $y^{\prime}$ and $z^{\prime}$, and to assume that $\Psi_{0}$ is a function of the single dimensionless variable $X=k_{1}x^{\prime}$. This simplification is especially useful because the optimal wave vector $k_{1}$ depends on the Raman coupling $\Omega_{R}$, which in turn can cause the antiperiod of $\Psi_{0}$, $\pi/k_{1}$, to vary at each perturbative order (although, by symmetry, only even order corrections in $\Omega_{R}$ are permitted; see Sec. IV.3). After this change of variable, the antiperiodicity condition becomes $\Psi_{0}(X+\pi)=-\Psi_{0}(X)$, fixing the antiperiod at $\pi$ for all orders. Therefore, $\Psi_{0}$ is $2\pi$-periodic in $X$.
With all the above considerations in place, we now insert the perturbative expansions [Eqs. (21), (22), (23), and (24)] into Eq. (8) and collect terms of equal order. At perturbative order $l\geq 1$, we obtain the recurrence relation
$\begin{split}&{}\left[E_{R}\left(-\mathrm{i}\nabla_{X}-\sigma_{z}\right)^{2}-G_{ss}s_{p}\left(-\mathrm{i}\nabla_{X}-\sigma_{z}\right)+\mathcal{L}_{D}\right]\Psi_{0}^{(l)}\\ &{}+\mathcal{L}_{C}\Psi_{0}^{(l)*}=\left[\mu^{(l)}+\frac{\Delta\mu^{(l)}}{2}(-\mathrm{i}\nabla_{X})\right]\Psi_{0}^{(0)}-\mathcal{J}^{(l)}\,,\end{split}$ (45)
where $\nabla_{X}$ denotes derivation with respect to the dimensionless variable $X$. The matrices $\mathcal{L}_{\mathrm{D}}$ and $\mathcal{L}_{\mathrm{C}}$ are defined as
@@
$\mathcal{J}^{(l)}={}$ ${}-\frac{\hbar^{2}}{2m}\sum_{l_{2}=2}^{l}\sum_{l_{1}=1}^{l_{2}-1}k_{1}^{(l_{1})}k_{1}^{(l_{2}-l_{1})}\nabla_{X}^{2}\Psi_{0}^{(l-l_{2})}$
${}+\frac{\hbar^{2}k_{R}}{m}\sum_{l_{1}=1}^{l-1}k_{1}^{(l_{1})}\left(-\nabla_{X}^{2}+\mathrm{i}\sigma_{z}\nabla_{X}\right)\Psi_{0}^{(l-l_{1})}$
${}+\frac{\hbar\Omega_{R}}{2}\,\sigma_{x}\Psi_{0}^{(l-1)}$
${}-\sum_{l_{1}=1}^{l-1}\mu^{(l_{1})}\Psi_{0}^{(l-l_{1})}-\sum_{l_{1}=1}^{l-1}\frac{\Delta\mu^{(l_{1})}}{2}(-\mathrm{i}\nabla_{X})\Psi_{0}^{(l-l_{1})}$
@@
In deriving Eqs. (46), we have set $\theta=\Delta\theta=0$ in the zero coupling expression (18) for the stripe order parameter, without loss of generality. Equation (45) thus provides a recursive method for determining the $l$th order corrections to the order parameter $\Psi_{0}^{(l)}$, as well as to the average chemical potential $\mu^{(l)}$ and the chemical potential difference $\Delta\mu^{(l)}$, once all lower-order corrections [which enter the source term $\mathcal{J}^{(l)}$] are known. As in standard time-independent perturbation theory, the first step is to extract $\mu^{(l)}$ and $\Delta\mu^{(l)}$. To do this, we multiply both sides of Eq. (45) by $\Psi_{0}^{(0)\dagger}$ and integrate over the full period of the wave function (from $X=-\pi$ to $X=\pi$). We then repeat this procedure with the operator $\Psi_{0}^{(0)\dagger}(-\mathrm{i}\nabla_{X})$. This yields two linear equations for $\mu^{(l)}$ and $\Delta\mu^{(l)}$, that can be readily solved. While the resulting expressions typically depend on the yet unknown $\Psi_{0}^{(l)}$, this dependence appears only through two integrals, which can be expressed using lower order quantities. These are
@@
$\begin{split}&{}\int_{-\pi}^{\pi}\mathrm{d}X\left[\Psi_{0}^{(0)\dagger}(-\mathrm{i}\nabla_{X})\Psi_{0}^{(l)}+\Psi_{0}^{(l)\dagger}(-\mathrm{i}\nabla_{X})\Psi_{0}^{(0)}\right]\\ &{}=-\sum_{l_{1}=1}^{l-1}\int_{-\pi}^{\pi}\mathrm{d}X\,\Psi_{0}^{(l_{1})\dagger}(-\mathrm{i}\nabla_{X})\Psi_{0}^{(l-l_{1})}\,.\end{split}$ (49)
These two relations are obtained by imposing the conservation of particle number [Eq. (4)] and momentum [Eq. (5)] at each perturbative order.
@@
$\Psi_{\mathrm{Hom}}^{(l)}=\mathrm{i}\left[\theta^{(l)}+\Delta\theta^{(l)}(-\mathrm{i}\nabla_{X})\right]\Psi_{0}^{(0)}\,,$ (50)
@@
${}\int_{-\pi}^{\pi}\mathrm{d}X\,\operatorname{Im}\left[\Psi_{0}^{(0)\dagger}(-\mathrm{i}\nabla_{X})\Psi_{0}^{(l)}\right]=0\,.$ (51b)
The function $\Psi_{0}^{(l)}$ computed following the above prescriptions is a sum of plane-wave components whose wave vectors (in units of $k_{1}$) are all odd integers from $-(2l+1)$ to $12l+$ 1. Thus, the number of harmonics increases with the perturbative order, in agreement with the Bloch-wave structure of the exact order parameter, Eq. (14). In the next sections, we explicitly compute the corrections at first and second order in $\hbar\Omega_{R}/4E_{R}$.

### v2

@@ V.2 Bogoliubov spectrum and sound velocities
The occurrence of an energetic instability in the spectrum of moving stripes is consistent with the behavior observed on the plane-wave side of the phase diagram. In this regime, the roton gap progressively decreases as the Raman coupling $\Omega_{R}$ is lowered towards the transition to the stripe phase [69, 89, 90]. Eventually, the gap closes at a spinodal point slightly below the critical value $\Omega_{\mathrm{cr1}}$. Upon further reduction of $\Omega_{R}$, the Bogoliubov modes near the roton minimum acquire negative frequency. These effects can be seen in the dashed lines of Fig. 5(d1)-(d2), although the plane-wave spectrum is modified in the comoving frame due to the Doppler shift term $vk_{x}$. In contrast, the plane-wave phase remains dynamically stable for all values of $\Omega_{R}$. This implies that the dynamical instability of moving stripe patterns must vanish once the limit $s_{p}\to\pm 1$ is reached. Our numerical analysis confirms this expectation: beyond a certain threshold, the dynamically unstable region begins to shrink as $|s_{p}|$ increases, and it disappears entirely when $|s_{p}|=1$. This behavior reflects the trend observed in the sound velocities, as discussed below.
@@
Figure 6: Velocities of spin ($b=1$, red lines) and density ($b=2$, blue lines) sound waves as functions of $s_{p}$ for (a) $\hbar\Omega_{R}/E_{R}=1.0$ and (b) $\hbar\Omega_{R}/E_{R}=2.6$. Results for both forward ($c_{b,+}$) and backward ($c_{b,-}$) propagation relative to the $x$ direction are shown. Owing to the symmetry property $c_{b,+}(-s_{p})=c_{b,-}(s_{p})$, we restrict the plot to positive values of $s_{p}$. For the spin sound velocities, we plot both the real parts (solid lines) and the imaginary parts (dashed lines), with the latter being identical for $c_{1,+}$ and $c_{1,-}$. The interaction parameters are the same as in previous figures: $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$.
@@
We now analyze in more detail the long wavelength behavior of the two gapless excitation bands ($b=1,2$), focusing in particular on the velocities of sound waves propagating parallel ($c_{b,+}$) and antiparallel ($c_{b,-}$) to the $x$ direction. These velocities characterize the linear dispersions of the gapless bands near the center of the first Brillouin zone, namely, $\omega_{b,k_{x}}=c_{b,+}k_{x}+O(k_{x}^{3})$ for $k_{x}>0$ and $\omega_{b,k_{x}}=-c_{b,-}k_{x}+O(k_{x}^{3})$ for $k_{x}<0$. As illustrated in Fig. 6, for both the spin ($b=1$) and density ($b=2$) sound modes, the velocities in the positive $x$ direction differ from those in the negative $x$ direction, except at $s_{p}=0$, where the two become equal. This symmetry at $s_{p}=0$ reflects the underlying symmetry of the excitation spectrum with respect to the center of the Brillouin zone, as discussed earlier. More generally, at arbitrary $s_{p}$ one has $c_{b,+}(-s_{p})=c_{b,-}(s_{p})$.
@@
The density sound velocities $c_{2,\pm}$ are always real and positive. In the limit $|s_{p}|\to 1$, they approach the expressions $c_{\pm x}\mp v_{x}$, where $c_{+x}$ ($c_{-x}$) denotes the speed of density sound waves propagating along (against) the $x$ direction in the plane-wave phase. Note that $c_{+x}\neq c_{-x}$ in the presence of nonzero spin-dependent interactions ($g_{ss}\neq 0$[69].
@@
In contrast, the behavior of the spin sound velocities reveals the onset of instability. As $|s_{p}|$ increases, one of the two velocities $c_{1,\pm}$ becomes negative at a certain point, indicating the emergence of an energetic instability, as previously discussed. For high Raman coupling, this instability is soon followed by the appearance of complex sound velocities: both $c_{1,+}$ and $c_{1,-}$ acquire imaginary parts of equal magnitude. These imaginary parts first grow with $|s_{p}|$, then decrease, and vanish as $|s_{p}|\to 1$. In this limit, the real parts tend towards opposite values, satisfying $c_{1,+}=-c_{1,-}$. This behavior confirms that the emergence of energetic instability is a necessary precursor to ensuring continuity of the excitation spectrum slope in the comoving frame at the points $k_{x}=\pm 2k_{1}$ in the plane-wave phase.
VI Experimental perspectives
To discuss the feasibility of observing moving stripe patterns, we first note that the interaction parameters used in this work differ significantly from those of the original experiment reported in Ref. [70]. The states of 87Rb employed in that experiment are indeed characterized by a very low degree of miscibility, as evidenced by the small ratio $g_{ss}/g_{dd}\sim 10^{-3}$, resulting in a small critical Raman coupling $\hbar\Omega_{\mathrm{cr1}}/E_{R}=0.19$. This limits the maximum achievable contrast of the stripes and thus restricts the possibility of observing significant supersolidity effects. Several recent works [20, 21, 22] have implemented strategies to address this limitation, leading to the detection of the density modulations characterizing the stripe phase. For example, by implementing spin-orbit coupling between two orbital states within a superlattice potential [91], it has been possible to reduce the spatial overlap between the wave functions of the two bare spin components along a direction parallel to the stripes (a similar effect can also be achieved by applying a quasi-two-dimensional spin-dependent trapping potential [92, 93, 94]). This resulted in a significant increase in the effective value of the ratio $g_{ss}/g_{dd}$, of more than two orders of magnitude compared to Ref. [70], and led to the first observation of the stripes by Ketterle’s group at MIT [20]. Alternatively, one can use atomic species with tunable interactions, as demonstrated in [22] employing 41K atoms. Stripes with high contrast can also be obtained by rapidly ramping the Raman coupling to some very large value, as done in [21]. Although these strategies have mostly, though not exclusively (see below), been employed to detect stationary stripes, they can also be used to observe traveling stripes, for which operating at sufficiently large Raman coupling is crucial to compensate for the reduction in contrast due to the stripe motion.
Moving stripe patterns can also be observed in configurations with finite Raman detuning or asymmetric intraspecies couplings. In these cases, stationary stripes are characterized by a nonzero dressed spin imbalance $s_{p}^{\mathrm{st}}$; traveling stripes then appear for $s_{p}\neq s_{p}^{\mathrm{st}}$. A first demonstration of this mechanism was provided by the MIT experiment [20], which reported the emergence of moving stripes with $s_{p}=0$ following the application of a Raman detuning. This generates a chemical potential difference between the two dressed spins even when their populations are equal. In contrast, the configuration explored in this work involves a population imbalance between the dressed spins, with the resulting stripe motion driven by spin-dependent interactions. This scenario can be realized experimentally by first preparing the BEC in a superposition of two atomic hyperfine states with different weights. Raman coupling can then be adiabatically introduced and ramped up to the desired strength. In typical experimental conditions, such as those of Refs. [20, 22], the fringe velocity can reach values up to a few millimeters per second, which lies within current experimental resolution. The traveling density modulations can be probed using Bragg scattering: a moving fringe pattern imparts a phase shift to the diffracted beam, which can be detected through interferometric techniques.
VII Conclusions
@@
In this work, we have investigated the properties of traveling supersolid patterns in a spin-orbit-coupled Bose-Einstein condensate. The motion of the density fringes stems from a population imbalance between the two minima of the single-particle dispersion. Antiferromagnetic spin-dependent interactions create a chemical potential difference which sets the stripe velocity. These moving stripes exhibit asymmetric densities of the two bare spin components, notable bare spin polarization, reduced contrast, and higher energy than that of stationary stripes. As the dressed spin imbalance increases, these effects intensify, eventually leading to the disappearance of the fringes and the emergence of a uniform plane-wave condensate. This indicates that traveling stripes interpolate between the stationary stripe and plane-wave phases. By analyzing the excitation spectrum, we uncover energetic and dynamical instabilities at high dressed spin imbalance, mainly affecting the spin phonon branch that becomes the mode with the roton minimum in the plane-wave phase. Our findings extend prior work [67, 68] on the slow motion of stripes induced by the release of a weak spin perturbation, where the propagation of fringes manifests the crystal Goldstone mode of the stripe phase. The parameter regime required to observe traveling stripe patterns is within reach of current experimental setups, opening new perspectives for the exploration of dynamical aspects of supersolidity in ultracold quantum gases.
@@ Appendix A Perturbative method
This appendix presents a detailed methodology employed to derive the perturbative formulas discussed in the main text. The general procedure described in Sec. A.1 closely follows the approach of Ref. [60], with appropriate generalizations to incorporate the additional constraint on dressed spin polarization. The first and second order corrections to the order parameter are subsequently derived in Sec. A.2 and Sec. A.3, respectively.
@@ A.1 General outline of the perturbative approach
To perturbatively solve the time-independent Gross-Pitaevskii equation (18), we first perform a redefinition of the order parameter, $\Psi_{0}(x^{\prime})\to\mathrm{e}^{\mathrm{i}k_{c}x^{\prime}}\Psi_{0}(X)$, so that $\Psi_{0}$ is now a function of the single dimensionless variable $X=k_{1}x^{\prime}$ satisfying the antiperiodicity condition $\Psi_{0}(X+\pi)=-\Psi_{0}(X)$ ($\Psi_{0}$ is thus $2\pi$-periodic in $X$). With these variable changes Eq. (18) takes the form
${}\left[h_{X}+\frac{\hbar\Omega_{R}}{2}\sigma_{x}+g_{dd}\left(\Psi_{0}^{\dagger}\Psi_{0}\right)+g_{ss}\left(\Psi_{0}^{\dagger}\sigma_{z}\Psi_{0}\right)\sigma_{z}\right]\Psi_{0}$
${}=(\mu_{d}+\mu_{s}p_{X})\Psi_{0}\,$, (55)
where we have defined the operators $p_{X}=-\mathrm{i}\nabla_{X}$ and
$h_{X}=E_{R}\left(\alpha_{1}p_{X}-\sigma_{z}\right)^{2}+2E_{R}\alpha_{c}\left(\alpha_{1}p_{X}-\sigma_{z}\right)+E_{R}\alpha_{c}^{2}\,$, (56)
as well as the dimensionless wave vectors $\alpha_{c}=k_{c}/k_{R}$ and $\alpha_{1}=k_{1}/k_{R}$. At $\Omega_{R}=0$, the uniform solutions of Eq. (55) take the form (30), only with $k_{R}x^{\prime}$ replaced by $X$. For arbitrary $\alpha_{c}$ and $\alpha_{1}$, these solutions have average energy density equal to
[…]
and from the thermodynamic relations (23) one deduces
$\mu_{d}^{(0)}={}$ ${}E_{R}(\alpha_{1}-1)^{2}+E_{R}\alpha_{c}^{2}+G_{dd}\,$, (58a)
$\mu_{s}^{(0)}={}$ ${}2E_{R}\alpha_{c}(\alpha_{1}-1)+G_{ss}s_{p}\,.$ (58b)
Note that $\bar{\varepsilon}^{(0)}$ is minimized by the choice $\alpha_{c}=0$ and $\alpha_{1}=1$, which leads to the results of Sec. IV.1.
Let us now insert the expansions (27), (28), and (29) into Eq. (55) and collect terms of equal order in the small parameter $\hbar\Omega_{R}/4E_{R}$. At perturbative order $l\geq 1$, we obtain the recurrence relation
$\begin{split}&{}\Big[E_{R}\left(\alpha_{1}p_{X}-\sigma_{z}\right)^{2}+\left(2E_{R}\alpha_{c}-G_{ss}s_{p}\right)\left(p_{X}-\sigma_{z}\right)\\ &{}-E_{R}\left(\alpha_{1}-1\right)^{2}+\mathcal{L}_{D}\Big]\Psi_{0}^{(l)}+\mathcal{L}_{C}\Psi_{0}^{(l)*}\\ &{}=\left[\mu_{d}^{(l)}+\mu_{s}^{(l)}p_{X}\right]\Psi_{0}^{(0)}-\mathcal{J}^{(l)}\,.\end{split}$ (59)
The matrices $\mathcal{L}_{\mathrm{D}}$ and $\mathcal{L}_{\mathrm{C}}$ are defined as
@@
$\mathcal{J}^{(l)}={}$ ${}\frac{\hbar\Omega_{R}}{2}\,\sigma_{x}\Psi_{0}^{(l-1)}-\sum_{l_{1}=1}^{l-1}\left[\mu_{d}^{(l_{1})}+\mu_{s}^{(l_{1})}p_{X}\right]\Psi_{0}^{(l-l_{1})}$
@@
In deriving Eqs. (60), we have set $\theta=\Delta\theta=0$ in the expression (30) for the unperturbed order parameter, without loss of generality. Equation (59) provides a recursive method for determining the $l$th order correction to the order parameter, $\Psi_{0}^{(l)}$, once all lower-order corrections are known. The latter are required to evaluate $\mathcal{J}^{(l)}$ and the correction to the average energy density, $\bar{\varepsilon}^{(l)}$, from which one obtains $\mu_{d}^{(l)}$ and $\mu_{s}^{(l)}$ using again the relations (23). Specifically, $\bar{\varepsilon}^{(l)}$ can be computed inserting the expansion (27) into the energy (3) divided by $V$. After singling out the term of order $l$, changing integration variable from $x$ to $X$, and carrying out the trivial integration over $y$ and $z$, one gets an expressions that depends on the yet unknown $\Psi_{0}^{(l)}$. However, this dependence appears only through two integrals, which can be expressed using lower order quantities. These are
@@
$\begin{split}&{}\int_{-\pi}^{\pi}\mathrm{d}X\left[\Psi_{0}^{(0)\dagger}p_{X}\Psi_{0}^{(l)}+\Psi_{0}^{(l)\dagger}p_{X}\Psi_{0}^{(0)}\right]\\ &{}=-\sum_{l_{1}=1}^{l-1}\int_{-\pi}^{\pi}\mathrm{d}X\,\Psi_{0}^{(l_{1})\dagger}p_{X}\Psi_{0}^{(l-l_{1})}\,.\end{split}$ (63)
For simplicity, here we are restricting integration over a single period of the order parameter, from $X=-\pi$ to $X=\pi$. Equations (62) and (63) are obtained by imposing the conservation of particle number [Eq. (4)] and momentum [Eq. (5)] at each perturbative order. The relation $p_{X}\Psi_{0}^{(0)}=\sigma_{z}\Psi_{0}^{(0)}$ must also be employed in the calculation of $\bar{\varepsilon}^{(l)}$.
Aside from the evaluation of $\mu_{d}^{(l)}$ and $\mu_{s}^{(l)}$, the knowledge of all the energy density corrections up to order $l$ is also crucial to determine the optimal values of the dimensionless wave vectors $\alpha_{c}$ and $\alpha_{1}$ up to the same order. This is achieved by imposing the two stationarity conditions $\frac{\partial}{\partial\alpha_{c}}\sum_{l_{1}=0}^{l}\bar{\varepsilon}^{(l_{1})}=0$ and $\frac{\partial}{\partial\alpha_{1}}\sum_{l_{1}=0}^{l}\bar{\varepsilon}^{(l_{1})}=0$. We then seek solutions in the form of perturbative expansions, $\alpha_{c}=\sum_{l_{1}=0}^{l}\alpha_{c}^{(l_{1})}$ and $\alpha_{1}=\sum_{l_{1}=0}^{l}\alpha_{1}^{(l_{1})}$, where the unperturbed values $\alpha_{c}^{(0)}=0$ and $\alpha_{1}^{(0)}=1$ coincide with those obtained previously. This self-consistent procedure allows the corrections $\alpha_{c}^{(l)}$ and $\alpha_{1}^{(l)}$ to be determined iteratively, order by order.
@@
$\Psi_{\mathrm{Hom}}^{(l)}=\mathrm{i}\left[\theta^{(l)}+\Delta\theta^{(l)}p_{X}\right]\Psi_{0}^{(0)}\,$, (64)
@@
${}\int_{-\pi}^{\pi}\mathrm{d}X\,\operatorname{Im}\left[\Psi_{0}^{(0)\dagger}p_{X}\Psi_{0}^{(l)}\right]=0\,.$ (65b)
The function $\Psi_{0}^{(l)}$ computed following the above prescriptions is a sum of plane-wave components whose wave vectors (in units of $k_{1}$) are all odd integers from (2l+1)-(2l+1) to $2l+1$. Thus, the number of harmonics increases with the perturbative order, in agreement with the Bloch-wave structure of the exact order parameter, Eq. (19). In the next sections, we explicitly compute the corrections at first and second order in $\hbar\Omega_{R}/4E_{R}$.